### v3

Abstract
This note proposes a new reformulation of complex semidefinite programs (SDPs) as real SDPs. As an application, we present an economical reformulation of complex SDP relaxations of complex polynomial optimization problems as real SDPs and derive some further reductions by exploiting inner structure of the complex SDP relaxations. Various numerical examples demonstrate that our new reformulation runs significantly faster than the usual popular reformulation.
keywords:
complex semidefinite programming, complex polynomial optimization, semidefinite programming, the complex moment-HSOS hierarchy, quantum information
{MSCcodes}
90C22, 90C23
1 Introduction
Complex semidefinite programs (SDPs) arise from a diverse set of areas, such as combinatorial optimization [9], optimal power flow [10, 12], quantum information theory [2, 4, 17], signal processing [11, 14]. In particular, they appear as convex relaxations of complex polynomial optimization problems (CPOPs), giving rise to the complex moment-Hermitian-sum-of-squares (moment-HSOS) hierarchy [10, 15, 16]. However, most modern SDP solvers deal with only real SDPs [^1] As far as the author knows, SeDuMi [13], Sdolab [7], and Hypatia [6] are the only solvers that can handle complex SDPs directly.. In order to handle complex SDPs via real SDP solvers, it is mandatory to reformulate complex SDPs as equivalent real SDPs. A popular way [^2] See for instance the online modeling cookbook of the commercial SDP solver MOSEK: https://docs.mosek.com/modeling-cookbook/sdo.html. to do so is to use the equivalent condition
(1) $H\succeq 0\quad\iff\quad Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0$
for an Hermitian matrix variable $H=H_{R}+H_{I}\hbox{\bf{i}}\in{\mathbb{C}}^{n\times n}$ with $H_{R}$ and $H_{I}$ being its real and imaginary parts respectively. Note that the right-hand-side constraint in (1) entails certain structure and to feed it to an SDP solver, we need to impose extra affine constraints to the positive semidefinite (PSD) constraint $Y\succeq 0$:
(2) $Y_{i,j}=Y_{i+n,j+n},Y_{i,j+n}+Y_{j,i+n}=0,\quad i=1,\ldots,n,j=i,\ldots,n.$
This conversion is quite simple but could be inefficient when $n$ is large. In this note, inspired by Lagrange duality, we propose a new reformulation of complex SDPs as real SDPs. The benefit of this new reformulation is that there is no need to add extra affine constraints and hence it owns lower complexity. In the same manner, we can obtain a new reformulation of complex SDP relaxations of CPOPs as real SDPs. Moreover, by exploiting inner structure of the complex SDP relaxations, we are able to remove a bunch of redundant affine constraints, which leads to an even more economical real reformulation of the complex SDP relaxations. Various numerical experiments (including randomly generated CPOPs and the alternating current optimal power flow (AC-OPF) problem) confirm our theoretical expectation and demonstrate that the new reformulation is more efficient than the usual popular one. Actually, our implementation of the new reformulation with MOSEK [1] also runs much faster than the implementation of the original complex formulation with Hypatia [8], probably because the SDP solvers based on real numbers are more mature and robust.
[…]
We endow ${\mathbb{C}}^{m}$ (viewed as a ${\mathbb{R}}$-vector space) with the scalar product $\langle\cdot,\cdot\rangle_{{\mathbb{R}}}$ defined by
$\langle u,v\rangle_{{\mathbb{R}}}={\mathcal{R}}(u^{{\mathrm{H}}}v)={\mathcal{R}}(u)^{\intercal}{\mathcal{R}}(v)+{\mathcal{I}}(u)^{\intercal}{\mathcal{I}}(v),\quad u,v\in{\mathbb{C}}^{m}.$
For $u,v\in{\mathbb{R}}^{m}$, $\langle u,v\rangle\coloneqq u^{\intercal}v$, where $u^{\intercal}$ stands for the transpose of $u$. For each $A\in{\mathbb{C}}^{n\times n}$, we associate it with an Hermitian matrix ${\mathcal{H}}(A)\coloneqq\frac{1}{2}(A+A^{{\mathrm{H}}})$. One can check that ${\mathcal{R}}(\langle A,H\rangle)=\langle{\mathcal{H}}(A),H\rangle$ for any $H\in{\mathrm{H}}^{n}$.
2 The real reformulations of complex SDPs
Given a tuple of complex matrices $A_{1},\ldots,A_{m}\in{\mathbb{C}}^{n\times n}$, we define a ${\mathbb{R}}$-linear operator ${\mathscr{A}}\colon{\mathrm{H}}^{n}\rightarrow{\mathbb{C}}^{m}$ by
(3) ${\mathscr{A}}(H)\coloneqq(\langle A_{i},H\rangle)_{i=1}^{m}\in{\mathbb{C}}^{m},\quad\forall H\in{\mathrm{H}}^{n}.$
Let us consider the following complex SDP:
($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{H\in{\mathrm{H}}^{n}}&\langle C,H\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}(H)=b,\\ &H\succeq 0,\end{cases}$
where $C\in{\mathrm{H}}^{n},b\in{\mathbb{C}}^{m}$. In order to convert ($-{\mathbb{C}}$) to a real SDP, we define two real linear operators ${\mathscr{A}}_{R},{\mathscr{A}}_{I}\colon{\mathbb{R}}^{n\times n}\rightarrow{\mathbb{R}}^{m}$ associated to ${\mathscr{A}}$ by
(4) ${\mathscr{A}}_{R}(S)\coloneqq\left(\langle{\mathcal{R}}(A_{i}),S\rangle\right)_{i=1}^{m}\in{\mathbb{R}}^{m},\quad\forall S\in{\mathbb{R}}^{n\times n}$
and
(5) ${\mathscr{A}}_{I}(S)\coloneqq\left(\langle{\mathcal{I}}(A_{i}),S\rangle\right)_{i=1}^{m}\in{\mathbb{R}}^{m},\quad\forall S\in{\mathbb{R}}^{n\times n},$
respectively. Moreover, assume $H=H_{R}+H_{I}\hbox{\bf{i}},C=C_{R}+C_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$ with $H_{R},H_{I},C_{R},C_{I}\in{\mathbb{R}}^{n\times n}$, $b_{R},b_{I}\in{\mathbb{R}}^{m}$. We can now convert ($-{\mathbb{C}}$) to a real SDP:
($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{Y\in{\mathbf{S}}^{2n}}&\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(H_{R})+{\mathscr{A}}_{I}(H_{I})=b_{R},\\ &{\mathscr{A}}_{R}(H_{I})-{\mathscr{A}}_{I}(H_{R})=b_{I},\\ &Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0.\end{cases}$
As mentioned in the introduction, to feed the PSD constraint in ($-{\mathbb{R}}$) to an SDP solver, we need to include also the extra $n(n+1)$ affine constraints listed in (2), which could be inefficient in practice. Below we show that by taking a dual point of view, we can actually get rid of this issue.
Before formulating the dual problem of ($-{\mathbb{C}}$), we explicitly give the adjoint operator of ${\mathscr{A}}$.
Lemma 2.1.
The adjoint operator ${\mathscr{A}}^{*}$ of ${\mathscr{A}}$ satisfies ${\mathscr{A}}^{*}(y)={\mathcal{H}}\left(\sum_{i=1}^{m}y_{i}A_{i}\right)$ for $y\in{\mathbb{C}}^{m}$.
Proof 2.2.
For any $H\in{\mathrm{H}}^{n}$, we have
$\left\langle{\mathscr{A}}^{*}(y),H\right\rangle$ $=\left\langle y,{\mathscr{A}}(H)\right\rangle_{{\mathbb{R}}}={\mathcal{R}}\left(\sum_{i=1}^{m}\overline{y}_{i}\langle A_{i},H\rangle\right)$
$={\mathcal{R}}\left(\left\langle\sum_{i=1}^{m}y_{i}A_{i},H\right\rangle\right)=\left\langle{\mathcal{H}}\left(\sum_{i=1}^{m}y_{i}A_{i}\right),H\right\rangle,$
which yields ${\mathscr{A}}^{*}(y)={\mathcal{H}}\left(\sum_{i=1}^{m}y_{i}A_{i}\right)$.
Then by the convex duality theory, the dual problem of ($-{\mathbb{C}}$) reads as
($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y\in{\mathbb{C}}^{m}}&\langle b,y\rangle_{{\mathbb{R}}}\\ \,\,\,\rm{s.t.}&{\mathscr{A}}^{*}(y)\succeq C.\end{cases}$
Assume $y=y_{R}+y_{I}\hbox{\bf{i}}$ with $y_{R},y_{I}\in{\mathbb{R}}^{m}$. Using Lemma 2.1, we deduce that ${\mathscr{A}}^{*}(y)=U+V\hbox{\bf{i}}$ with
$U\coloneqq\frac{1}{2}\sum_{i=1}^{m}{\mathcal{R}}(y_{i}){\mathcal{R}}(A_{i}+A_{i}^{\intercal})-{\mathcal{I}}(y_{i}){\mathcal{I}}(A_{i}+A_{i}^{\intercal})$
and
$V\coloneqq\frac{1}{2}\sum_{i=1}^{m}{\mathcal{I}}(y_{i}){\mathcal{R}}(A_{i}-A_{i}^{\intercal})+{\mathcal{R}}(y_{i}){\mathcal{I}}(A_{i}-A_{i}^{\intercal}).$
Thus, we can convert ($-{\mathbb{C}}$) to a real SDP by using the equivalent condition (1):
($-{\mathbb{R}}$) $\begin{cases}\inf\limits_{y_{R},y_{I}\in{\mathbb{R}}^{m}}&b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}\\ \,\,\,\,\,\,\,\,\rm{s.t.}&\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\succeq 0.\end{cases}$
Let $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]\in{\mathbf{S}}^{2n}$ be the dual PSD variable of ($-{\mathbb{R}}$) with $X_{1},X_{2},X_{3}\in{\mathbb{R}}^{n\times n}$. Then the Lagrangian associated with ($-{\mathbb{R}}$) given by
$\,L(X,y_{R},y_{I})$
$=$ $\,b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}-\left\langle\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix},\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\right\rangle$
$=$ $\,b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}-\langle X_{1}+X_{2},U-C_{R}\rangle-\langle X_{3}-X_{3}^{\intercal},V-C_{I}\rangle$
$=$ $\,b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}+\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$
$\,-\sum_{i=1}^{m}\langle{\mathcal{R}}(A_{i}),X_{1}+X_{2}\rangle{\mathcal{R}}(y_{i})+\sum_{i=1}^{m}\langle{\mathcal{I}}(A_{i}),X_{1}+X_{2}\rangle{\mathcal{I}}(y_{i})$
$\,-\sum_{i=1}^{m}\langle{\mathcal{R}}(A_{i}),X_{3}-X_{3}^{\intercal}\rangle{\mathcal{I}}(y_{i})-\sum_{i=1}^{m}\langle{\mathcal{I}}(A_{i}),X_{3}-X_{3}^{\intercal}\rangle{\mathcal{R}}(y_{i})$
$=$ $\,\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle+\langle b_{R}-{\mathscr{A}}_{R}(X_{1}+X_{2})-{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal}),y_{R}\rangle$
$\,+\langle b_{I}-{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})+{\mathscr{A}}_{I}(X_{1}+X_{2}),y_{I}\rangle.$
Therefore, the dual problem of ($-{\mathbb{R}}$) can be written down as
($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{X\in{\mathbf{S}}^{2n}}&\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(X_{1}+X_{2})+{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal})=b_{R},\\ &{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})-{\mathscr{A}}_{I}(X_{1}+X_{2})=b_{I},\\ &X=\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix}\succeq 0.\end{cases}$
Theorem 2.3.
($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{R}}$) (in the sense that they share the same optimum). As a result, ($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{C}}$). In addition, if $X^{\star}=\left[\begin{smallmatrix}X_{1}^{\star}&(X_{3}^{\star})^{\intercal}\\ X_{3}^{\star}&X_{2}^{\star}\end{smallmatrix}\right]$ is an optimal solution to ($-{\mathbb{R}}$), then $H^{\star}=(X^{\star}_{1}+X_{2}^{\star})+(X^{\star}_{3}-(X_{3}^{\star})^{\intercal})\hbox{\bf{i}}$ is an optimal solution to ($-{\mathbb{C}}$).
Proof 2.4.
Let us denote the optima of ($-{\mathbb{R}}$) and ($-{\mathbb{R}}$) by $v$ and $v^{\prime}$, respectively. Suppose that $Y=\left[\begin{smallmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$). Then one can easily check that $X\coloneqq\frac{1}{2}Y$ is a feasible solution to ($-{\mathbb{R}}$). Moreover, we have $\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle=\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle$ and it follows $v\leq v^{\prime}$. On the other hand, suppose $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$). We then have
$\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}^{-1}X\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}=\begin{bmatrix}X_{2}&-X_{3}\\ -X_{3}^{\intercal}&X_{1}\end{bmatrix}\succeq 0,$
and thus
$Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\coloneqq\begin{bmatrix}X_{1}+X_{2}&X_{3}^{\intercal}-X_{3}\\ X_{3}-X_{3}^{\intercal}&X_{1}+X_{2}\end{bmatrix}\succeq 0.$
One can easily see that $Y$ is a feasible solution to ($-{\mathbb{R}}$) and in addition, it holds $\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle=\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$. Thus $v\geq v^{\prime}$, which proves the equivalence. The latter statement of the theorem is clear from the above arguments.
In contrast to ($-{\mathbb{R}}$), the PSD constraint in ($-{\mathbb{R}}$) is straightforward, and thus no extra affine constraint is required. This is why the conversion ($-{\mathbb{R}}$) is more appealing than ($-{\mathbb{R}}$) from the computational perspective.
Remark 2.5.
A similar reformulation to ($-{\mathbb{R}}$) but for a restricted class of complex SDP relaxations of multiple-input multiple-output detection has appeared in [11].
3 Application to complex SDP relaxations for CPOPs
In this section, we apply the reformulation ($-{\mathbb{R}}$) to complex SDP relaxations arising from the complex moment-HSOS hierarchy for CPOPs. A CPOP is given by
(CPOP) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&f({\mathbf{z}},\overline{{\mathbf{z}}})=\sum_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}\\ \,\,\rm{s.t.}&g_{i}({\mathbf{z}},\overline{{\mathbf{z}}})=\sum_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}\geq 0,\quad i\in[t],\end{cases}$
where $\overline{{\mathbf{z}}}\coloneqq(\overline{z}_{1},\ldots,\overline{z}_{s})$ stands for the conjugate of complex variables ${\mathbf{z}}\coloneqq(z_{1},\ldots,z_{s})$. The functions $f,g_{1},\ldots,g_{t}$ are real-valued polynomials and their coefficients satisfy $b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$, $g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{g^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$. The support of $f$ is defined by $\hbox{\rm{supp}}(f)\coloneqq\{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\mid b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$. For $i\in[t]$, $\hbox{\rm{supp}}(g_{i})$ is defined in the same way.
Fix a $d\in{\mathbb{N}}$. Let $y=(y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}})_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}}\subseteq{\mathbb{C}}$ be a sequence indexed by $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}$ and satisfying $y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{y_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$. Let $L_{y}$ be the linear functional defined by
$f=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})}b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}\mapsto L_{y}(f)=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})}b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}.$
The complex moment matrix ${\mathbf{M}}_{d}(y)$ associated with $y$ is the Hermitian matrix indexed by ${\mathbb{N}}^{s}_{d}$ such that
$[{\mathbf{M}}_{d}(y)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}({\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
Suppose that $g=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}{\mathbf{z}}^{{\boldsymbol{\beta}}^{\prime}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}^{\prime}}$ is a complex polynomial. The complex localizing matrix ${\mathbf{M}}_{d}(gy)$ associated with $g$ and $y$ is the Hermitian matrix indexed by ${\mathbb{N}}^{s}_{d}$ such that
$[{\mathbf{M}}_{d}(gy)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}(g{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}y_{{\boldsymbol{\beta}}+{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}+{\boldsymbol{\gamma}}^{\prime}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
Let $d_{0}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ and $d_{i}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ for $i\in[t]$, where ||$\cdot$|| ⋅ | denotes the sum of entries. Let $d_{\min}\coloneqq\max\,\{d_{0},d_{1},\ldots,d_{m}\}$. For any $d\geq d_{\min}$, the $d$-th ($d$ is called the relaxation order) complex moment relaxation for (CPOP) is given by
($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y}&L_{y}(f)=\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&{\mathbf{M}}_{d}(y)\succeq 0,\\ &{\mathbf{M}}_{d-d_{i}}(g_{i}y)\succeq 0,\quad i\in[t],\\ &y_{\mathbf{0},\mathbf{0}}=1.\end{cases}$
($-{\mathbb{C}}$) and its dual problem form the complex moment-HSOS hierarchy of (CPOP). For more details on this hierarchy, we refer the reader to [10, 15].
For any $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d}$, we associate it with a matrix $A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{\omega_{s,d}\times\omega_{s,d}}$ defined by
(6) $[A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}1,&\text{if }({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
Moreover, for each $i\in[t]$, we associate any $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}$ with a matrix $A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{C}}^{\omega_{s,d-d_{i}}\times\omega_{s,d-d_{i}}}$ defined by
(7) $[A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}},&\text{if }({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
Now for each $i=0,1,\ldots,t$, we define the ${\mathbb{R}}$-linear operator ${\mathscr{A}}^{i}$ by
${\mathscr{A}}^{i}(H)\coloneqq\left(\langle A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},H\rangle\right)_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}},\quad H\in{\mathrm{H}}^{\omega_{s,d-d_{i}}}.$
For convenience let us set $g_{0}\coloneqq 1$. By construction, it holds
${\mathbf{M}}_{d-d_{i}}(g_{i}y)=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}}A_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}^{i}y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=({\mathscr{A}}^{i})^{*}(y),\quad i=0,1,\ldots,t.$
Therefore, one can rewrite ($-{\mathbb{C}}$) as follows:
($-{\mathbb{C}}$’) $\begin{cases}\inf\limits_{y}&\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&({\mathscr{A}}^{i})^{*}(y)\succeq 0,\quad i=0,1,\ldots,t,\\ &y_{\mathbf{0},\mathbf{0}}=1,\end{cases}$
whose dual reads as
($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Note that we have used the Kronecker delta $\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}$ in ($-{\mathbb{C}}$).
Let us from now on fix any order “$<$” on ${\mathbb{N}}^{s}$.
Proposition 3.1.
($-{\mathbb{C}}$) is equivalent to the following complex SDP:
($-{\mathbb{C}}$’) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Proof 3.2.
It suffices to show that for ${\boldsymbol{\beta}}<{\boldsymbol{\gamma}}$, $\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}=b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}$ is equivalent to $\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}$. Indeed, this equivalence follows from $b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}=\overline{b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}}$ and
$\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}$ $=\sum_{i=0}^{t}\langle A^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}},H^{i}\rangle$
$=[H^{0}]_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}+\sum_{i=1}^{t}\sum_{\begin{subarray}{c}({\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime})=({\boldsymbol{\gamma}},{\boldsymbol{\beta}})\end{subarray}}g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}}[H^{i}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$
$=\sum_{i=0}^{t}\langle\overline{A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}},H^{i}\rangle=\sum_{i=0}^{t}\overline{[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}}.$
Let ${\mathscr{A}}^{i}_{R},{\mathscr{A}}^{i}_{I}$ be the real linear operator associated with ${\mathscr{A}}^{i}$ which are defined in a similar way as (4) and (5). With $H^{i}=H^{i}_{R}+H^{i}_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$, ($-{\mathbb{C}}$) is equivalent to the following real SDP by using the equivalent condition (1):
($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{\lambda,Y^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &Y^{i}=\begin{bmatrix}H^{i}_{R}&-H^{i}_{I}\\ H^{i}_{I}&H^{i}_{R}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
On the other hand, by invoking Theorem 2.3, we obtain another equivalent real SDP conversion of ($-{\mathbb{C}}$):
(8) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Proposition 3.3.
(8) is equivalent to the following real SDP:
($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad{\boldsymbol{\beta}}\neq{\boldsymbol{\gamma}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Proof 3.4.
We need to show that the following constraints
(9) $\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=0,\quad{\boldsymbol{\beta}}\in{\mathbb{N}}^{s}_{d}$
in (8) are redundant. For each $i=0,1,\ldots,t$ and ${\boldsymbol{\beta}}\in{\mathbb{N}}^{s}_{d}$, we have
$\,\left\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\right\rangle$
$=$ $\,\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{R}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})\left([X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}-[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}\right)$
$=$ $\,\frac{1}{2}\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{R}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})\left([X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}+[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}-[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}-[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}\right)$
[…]
where we have used fact that $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$ and $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}$. It follows that $[{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\rangle=0$.
In addition, for each $i=0,1,\ldots,t$ and ${\boldsymbol{\beta}}\in{\mathbb{N}}^{s}_{d}$, we have
$\,\left\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\right\rangle$
$=$ $\,\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})\left([X^{i}_{1}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}+[X^{i}_{2}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}\right)$
$=$ $\,\frac{1}{2}\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{I}}\left(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}}\right)\left([X^{i}_{1}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}+[X^{i}_{2}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}\right)$
[…]
where we have used fact that ${\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}})={\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+\overline{g}^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})=0$ and $X^{i}_{1},X^{i}_{2}$ are symmetric. It follows that $[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\rangle=0$.
Putting all above together yields (9).
We have proved the following theorem.
Before closing the section, we compare complexity of different real SDP reformulations for complex SDP relaxations of (CPOP) in Table 1.
Table 1: Comparison of complexity of different real SDP reformulations for complex SDP relaxations of (CPOP). $n_{\rm{sdp}}$: the maximal size of SDP matrix, $m_{\rm{sdp}}$: the number of affine constraints.
($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
$n_{\rm{sdp}}$ $2\omega_{s,d}$ $2\omega_{s,d}$
$m_{\rm{sdp}}$ $2\omega_{s,d}^{2}+2\omega_{s,d}+\sum_{i=1}^{t}\omega_{s,d-d_{i}}$ $\omega_{s,d}^{2}$
4 Numerical experiments
In this section, we benchmark the performance of the two real reformulations for complex SDPs using the software TSSOS 1.2.1333TSSOS is freely available at https://github.com/wangjie212/TSSOS. in which MOSEK 10.0 [1] is employed as an SDP solver with default settings. For comparison, we also include the results of directly solving complex SDPs obtained with Hypatia 0.8.1 [6] in Sections 4.14.3. All numerical experiments were performed on a desktop computer with Intel(R) Core(TM) i9-10900 CPU@2.80GHz and 64G RAM. When presenting the results, the column labelled by ‘opt’ records the optimum and the column labelled by ‘time’ records running time in seconds. Moreover, the symbol ‘-’ means the SDP solver runs out of memory, and the symbol ‘*’ means running time exceeds 10000 seconds.
4.1 Minimizing a random complex quartic polynomial over the unit sphere
Our first example is to minimize a complex quartic polynomial over the unit sphere:
(10) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&[{\mathbf{z}}]_{2}^{{\mathrm{H}}}Q[{\mathbf{z}}]_{2}\\ \,\,\,\rm{s.t.}&|z_{1}|^{2}+\cdots+|z_{s}|^{2}=1,\end{cases}$
where $[{\mathbf{z}}]_{2}$ is the column vector of monomials in ${\mathbf{z}}$ up to degree two and $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the standard normal distribution.
We approach (10) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 2. From the table, we see that the reformulation ($-{\mathbb{R}}$) is several ($2\sim 7$) times as fast as the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows. On the other hand, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
Table 2: Minimizing a random complex quartic polynomial over the unit sphere.
$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
5 2 966 -11.2409 0.11 441 -11.2409 0.05 -11.2409 0.12
3 6846 -9.47725 8.13 3136 -9.47725 2.00 -9.47725 6.54
7 2 2736 -14.2314 0.97 1296 -14.2314 0.28 -14.2314 0.59
3 30372 -11.0407 389 14400 -11.0407 57.0 -11.0407 474
9 2 6270 -19.0019 5.73 3025 -19.0019 1.62 -19.0019 4.61
3 100320 - - 48400 -15.5614 1944 - -
11 2 12480 -22.8630 31.7 6084 -22.8630 6.67 -22.8630 32.3
3 271882 - - 132496 - - - -
13 2 22470 -25.6352 145 11025 -25.6352 23.5 -25.6352 174
3 639450 - - 313600 - - - -
15 2 37536 -29.1672 585 18496 -29.1672 86.1 -29.1672 802
3 1351976 - - 665856 - - - -
4.2 Minimizing a random complex quartic polynomial with unit-norm variables
The second example is to minimize a random complex quartic polynomial with unit-norm variables:
(11) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&[{\mathbf{z}}]_{2}^{{\mathrm{H}}}Q[{\mathbf{z}}]_{2}\\ \,\,\,\rm{s.t.}&|z_{i}|^{2}=1,\quad i=1,\ldots,s,\end{cases}$
where $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the uniform probability distribution on $[0,1]$.
We approach (11) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 3. From the table, we see that the reformulation ($-{\mathbb{R}}$) is about one magnitude faster than the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows. Again, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
Table 3: Minimizing a random complex quartic polynomial with unit-norm variables.
$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
5 2 734 -24.4919 0.10 271 -24.4919 0.03 -24.4919 0.21
3 4474 -24.4919 2.34 1281 -24.4919 0.26 -24.4919 10.9
7 2 2202 -56.5289 0.65 869 -56.5289 0.16 -56.5289 1.15
3 21158 -46.7128 132 6637 -46.7128 7.44 -46.7128 520
9 2 5242 -114.342 4.62 2161 -114.342 0.73 -114.342 5.29
3 73312 - - 24691 -81.2676 184 - -
11 2 10718 -202.436 32.1 4555 -202.436 3.86 -202.436 30.0
3 206188 - - 73327 - - - -
13 2 19686 -338.041 126 8555 -338.041 12.7 -338.041 162
3 499438 - - 185277 - - - -
15 2 33394 -514.226 678 14761 -514.226 55.1 -514.226 705
3 1081514 - - 414841 - - - -
4.3 Minimizing a randomly generated sparse complex quartic polynomial over multi-spheres
[…]
(12) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{5(l+1)}}&f({\mathbf{z}},\overline{{\mathbf{z}}})\\ \,\quad\rm{s.t.}&\sum_{j=1}^{10}|z_{5(i-1)+j}|^{2}=1,\quad i=1,\ldots,l.\end{cases}$
The sparsity in (12) can be exploited to derive a sparsity-adapted complex moment-HSOS hierarchy [15]. We solve the second sparse HSOS relaxation of (12) for $l=40,80,\ldots,400$. The results are displayed in Table 4. From the table we see that the reformulation ($-{\mathbb{R}}$) is $1.5\sim 2$ times as fast as the reformulation ($-{\mathbb{R}}$). Moreover, for this problem, solving the original complex SDP with Hypatia is extremely slow probably due to the fact that the SDP contains many PSD blocks.
Table 4: Minimizing a randomly generated sparse complex quartic polynomial over multi-spheres.
$l$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
40 23090 -98.9240 3.12 12529 -98.9240 2.06 -98.9240 886
80 46768 -197.577 12.6 25549 -197.577 8.07 -197.577 5433
120 70958 -292.024 30.1 38871 -292.024 19.0 * *
160 94278 -389.652 45.9 51563 -389.652 30.7 * *
200 117526 -482.684 84.5 64185 -482.684 37.7 * *
240 140298 -578.896 130 76389 -578.896 59.5 * *
280 162504 -671.047 173 89241 -671.047 65.4 * *
320 187528 -766.403 206 102171 -766.403 88.5 * *
360 210370 -866.771 291 114589 -866.771 147 * *
400 233396 -963.137 297 127173 -963.137 138 * *
4.4 Application to the AC-OPF problem
The AC-OPF is a central problem in power systems, which aims to minimize the generation cost of an alternating current transmission network under physical and operational constraints. Mathematically, it can be formulated as the following CPOP:
(13) $\begin{cases}\inf\limits_{V_{i},S_{k}^{g}}&\sum_{k\in G}\left(\mathbf{c}_{2k}({\mathcal{R}}(S_{k}^{g}))^{2}+\mathbf{c}_{1k}{\mathcal{R}}(S_{k}^{g})+\mathbf{c}_{0k}\right)\\ \,\,\,\textrm{s.t.}&\angle V_{r}=0,\\ &\mathbf{S}_{k}^{gl}\leq S_{k}^{g}\leq\mathbf{S}_{k}^{gu},\quad\forall k\in G,\\ &\boldsymbol{\upsilon}_{i}^{l}\leq|V_{i}|\leq\boldsymbol{\upsilon}_{i}^{u},\quad\forall i\in N,\\ &\sum_{k\in G_{i}}S_{k}^{g}-\mathbf{S}_{i}^{d}-\mathbf{Y}_{i}^{sh}|V_{i}|^{2}=\sum_{(i,j)\in E_{i}\cup E_{i}^{R}}S_{ij},\quad\forall i\in N,\\ &S_{ij}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})\frac{|V_{i}|^{2}}{|\mathbf{T}_{ij}|^{2}}-\overline{\mathbf{Y}}_{ij}\frac{V_{i}\overline{V}_{j}}{\mathbf{T}_{ij}},\quad\forall(i,j)\in E,\\ &S_{ji}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})|V_{j}|^{2}-\overline{\mathbf{Y}}_{ij}\frac{\overline{V}_{i}V_{j}}{\overline{\mathbf{T}}_{ij}},\quad\forall(i,j)\in E,\\ &|S_{ij}|\leq\mathbf{s}_{ij}^{u},\quad\forall(i,j)\in E\cup E^{R},\\ &\boldsymbol{\theta}_{ij}^{\Delta l}\leq\angle(V_{i}\overline{V}_{j})\leq\boldsymbol{\theta}_{ij}^{\Delta u},\quad\forall(i,j)\in E,\\ \end{cases}$
[…]
We select test cases from the AC-OPF library PGLiB-OPF [3]. For each case, we solve the minimal relaxation step of the sparse HSOS hierarchy [15]. The results are displayed in Table 5. From the table, we again see that the reformulation ($-{\mathbb{R}}$) is several ($1.4\sim 5$) times as fast as the reformulation ($-{\mathbb{R}}$).
Table 5: The results for the AC-OPF problem. $s$: the number of CPOP variables; $t$: the number of CPOP constraints.
Case $s$ $t$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time
[…]
Acknowledgments
The authors would like to thank Jurij Volčič for helpful comments on an earlier preprint of this note.

### v4

@@
-
+Jie Wang Academy of Mathematics and Systems Science, Chinese Academy of Sciences, Beijing, China (, https://wangjie212.github.io/jiewang/)
 Abstract
-This note proposes a new reformulation of complex semidefinite programs (SDPs) as real SDPs.
+This note proposes a new reformulation of complex semidefinite programs (SDP) as real SDPs.
 As an application, we present an economical reformulation of complex SDP relaxations of complex polynomial optimization problems as real SDPs and derive some further reductions by exploiting inner structure of the complex SDP relaxations.
@@
 1 Introduction
-Complex semidefinite programs (SDPs) arise from a diverse set of areas, such as combinatorial optimization [9], optimal power flow [10, 12], quantum information theory [2, 4, 17], signal processing [11, 14].
-In particular, they appear as convex relaxations of complex polynomial optimization problems (CPOPs), giving rise to the complex moment-Hermitian-sum-of-squares (moment-HSOS) hierarchy [10, 15, 16].
+Complex semidefinite programs (SDP) arise from a diverse set of areas, such as combinatorial optimization [11], optimal power flow [12, 14], quantum information theory [2, 4, 19], and signal processing [13, 16].
+In particular, they appear as convex relaxations of complex polynomial optimization problems (CPOP), giving rise to the complex moment-Hermitian-sum-of-squares (moment-HSOS) hierarchy [12, 17, 18].
 However, most modern SDP solvers deal with only real SDPs [^1] As far as the author knows, SeDuMi [15], Sdolab [9], and Hypatia [8] are the only solvers that can handle complex SDPs directly..
@@
 (1) $H\succeq 0\quad\iff\quad Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0$
-for an Hermitian matrix variable $H=H_{R}+H_{I}\hbox{\bf{i}}\in{\mathbb{C}}^{n\times n}$ with $H_{R}$ and $H_{I}$ being its real and imaginary parts respectively.
+for a Hermitian matrix variable $H=H_{R}+H_{I}\hbox{\bf{i}}\in{\mathbb{C}}^{n\times n}$ with $H_{R}$ and $H_{I}$ being its real and imaginary parts respectively.
 Note that the right-hand-side constraint in (1) entails certain structure and to feed it to an SDP solver, we need to impose extra affine constraints to the positive semidefinite (PSD) constraint $Y\succeq 0$:
@@
 In this note, inspired by Lagrange duality, we propose a new reformulation of complex SDPs as real SDPs.
-The benefit of this new reformulation is that there is no need to add extra affine constraints and hence it owns lower complexity.
+The benefit of this new reformulation is that it contains no extra affine constraints while keeping the same matrix size and hence has lower complexity.
 In the same manner, we can obtain a new reformulation of complex SDP relaxations of CPOPs as real SDPs.
 Moreover, by exploiting inner structure of the complex SDP relaxations, we are able to remove a bunch of redundant affine constraints, which leads to an even more economical real reformulation of the complex SDP relaxations.
-Various numerical experiments (including randomly generated CPOPs and the alternating current optimal power flow (AC-OPF) problem) confirm our theoretical expectation and demonstrate that the new reformulation is more efficient than the usual popular one.
-Actually, our implementation of the new reformulation with MOSEK [1] also runs much faster than the implementation of the original complex formulation with Hypatia [8], probably because the SDP solvers based on real numbers are more mature and robust.
-[…]
+Various numerical experiments (including the complex matrix completion problem, complex SDP relaxations for randomly generated CPOPs, and the alternating current optimal power flow (AC-OPF) problem) confirm our theoretical expectation and demonstrate that the new reformulation is substantially more efficient than the usual popular one.
+Finally, numerical experiments also indicate that our implementation of the new reformulation with MOSEK [1] runs much faster than the implementation of the original complex formulation with Hypatia [10], probably because MOSEK is commercial and the implementations of SDP algorithms based on real numbers are more mature and robust.
+Notation.
+The symbol $\mathbb{N}$} denotes the set of nonnegative integers.
+For $n\in{\mathbb{N}}\setminus\{0\}$, let $n]\coloneqq\{1,2,\ldots,n\}$.
+We write $|A|$ for the cardinality of a set $A$.
+Let i be the imaginary unit, satisfying $\hbox{\bf{i}}^{2}=-1$.
+For $d\in{\mathbb{N}}$, let ${\mathbb{N}}^{n}_{d}\coloneqq\{(\alpha_{i})_{i}\in{\mathbb{N}}^{n}\mid\sum_{i=1}^{n}\alpha_{i}\leq d\}$ and let $\omega_{n,d}\coloneqq\binom{n+d}{d}$ be the cardinality of ${\mathbb{N}}^{n}_{d}$.
+For ${\boldsymbol{\alpha}}=(\alpha_{i})_{i}\in{\mathbb{N}}^{n}_{d}$ and an $n$-tuple of variables ${\mathbf{z}}=\{z_{1},\ldots,z_{n}\}$, let ${\mathbf{z}}^{{\boldsymbol{\alpha}}}\coloneqq z_{1}^{\alpha_{1}}\cdots z_{n}^{\alpha_{n}}$.
+For a complex number $a$, $\overline{a}$ (resp.
+${\mathcal{R}}(a),{\mathcal{I}}(a)$) denotes the conjugate (resp.
+real part, imaginary part) of $a$, and for a complex vector $v$, $v^{{\mathrm{H}}}$ denotes the conjugate transpose of $v$.
+For a positive integer $n$, the set of $n\times n$ symmetric (resp.
+Hermitian) matrices is denoted by ${\mathbf{S}}^{n}$ (resp.
+${\mathrm{H}}^{n}$).
+We use $A\succeq 0$ to indicate that the matrix $A$ is PSD.
+For $A,B\in{\mathbb{C}}^{n\times n}$, we denote by $\langle A,B\rangle$ the trace inner-product, defined by $\langle A,B\rangle=\hbox{\rm{Tr}}(A^{{\mathrm{H}}}B)$, where $A^{{\mathrm{H}}}$ stands for the conjugate transpose of $A$.
+For $\in{\mathbb{R}}^{n\times n}, AA^{\intercal}$ stands for the transpose of $A$.
 We endow ${\mathbb{C}}^{m}$ (viewed as a $\mathbb{R}$}-vector space) with the scalar product $\langle\cdot,\cdot\rangle_{{\mathbb{R}}}$ defined by
-$\langle u,v\rangle_{{\mathbb{R}}}={\mathcal{R}}(u^{{\mathrm{H}}}v)={\mathcal{R}}(u)^{\intercal}{\mathcal{R}}(v)+{\mathcal{I}}(u)^{\intercal}{\mathcal{I}}(v),\quad u,v\in{\mathbb{C}}^{m}.$
-For $u,v\in{\mathbb{R}}^{m}$, $\langle u,v\rangle\coloneqq u^{\intercal}v$, where $u^{\intercal}$ stands for the transpose of $u$.
-For each $A\in{\mathbb{C}}^{n\times n}$, we associate it with an Hermitian matrix ${\mathcal{H}}(A)\coloneqq\frac{1}{2}(A+A^{{\mathrm{H}}})$.
+$\langle u,v\rangle_{{\mathbb{R}}}={\mathcal{R}}(u^{{\mathrm{H}}}v)={\mathcal{R}}(u)^{\intercal}{\mathcal{R}}(v)+{\mathcal{I}}(u)^{\intercal}{\mathcal{I}}(v),\quad u,v\in{\mathbb{C}}^{m},$
+which will be used in forming the dual complex SDP problem (to obtain a real objective).
+For $\in{\mathbb{R}}^{m}, u,vuv\langle$ $\rangle\coloneqq u^{\intercal}v$, where $u^{\intercal}$ stands for the transpose of $u$.
+We associate each $A\in{\mathbb{C}}^{n\times n}$ with a Hermitian matrix ${\mathcal{H}}(A)\coloneqq\frac{1}{2}(A+A^{{\mathrm{H}}})$.
 One can check that ${\mathcal{R}}(\langle A,H\rangle)=\langle{\mathcal{H}}(A),H\rangle$ for any $H\in{\mathrm{H}}^{n}$.
@@
 Let us consider the following complex SDP:
-($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{H\in{\mathrm{H}}^{n}}&\langle C,H\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}(H)=b,\\ &H\succeq 0,\end{cases}$
+(PSDP-$\mathbb{C}$}) $\begin{cases}\sup\limits_{H\in{\mathrm{H}}^{n}}&\langle C,H\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}(H)=b,\\ &H\succeq 0,\end{cases}$
 where $C\in{\mathrm{H}}^{n},b\in{\mathbb{C}}^{m}$.
-In order to convert ($-{\mathbb{C}}$) to a real SDP, we define two real linear operators ${\mathscr{A}}_{R},{\mathscr{A}}_{I}\colon{\mathbb{R}}^{n\times n}\rightarrow{\mathbb{R}}^{m}$ associated to ${\mathscr{A}}$ by
+In order to convert (PSDP-$\mathbb{C}$}) to a real SDP, we define two real linear operators ${\mathscr{A}}_{R},{\mathscr{A}}_{I}\colon{\mathbb{R}}^{n\times n}\rightarrow{\mathbb{R}}^{m}$ associated with $\mathscr{A}$} by
 (4) ${\mathscr{A}}_{R}(S)\coloneqq\left(\langle{\mathcal{R}}(A_{i}),S\rangle\right)_{i=1}^{m}\in{\mathbb{R}}^{m},\quad\forall S\in{\mathbb{R}}^{n\times n}$
@@
 Moreover, assume $H=H_{R}+H_{I}\hbox{\bf{i}},C=C_{R}+C_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$ with $H_{R},H_{I},C_{R},C_{I}\in{\mathbb{R}}^{n\times n}$, $b_{R},b_{I}\in{\mathbb{R}}^{m}$.
-We can now convert ($-{\mathbb{C}}$) to a real SDP:
-($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{Y\in{\mathbf{S}}^{2n}}&\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(H_{R})+{\mathscr{A}}_{I}(H_{I})=b_{R},\\ &{\mathscr{A}}_{R}(H_{I})-{\mathscr{A}}_{I}(H_{R})=b_{I},\\ &Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0.\end{cases}$
-As mentioned in the introduction, to feed the PSD constraint in ($-{\mathbb{R}}$) to an SDP solver, we need to include also the extra $n(n+1)$ affine constraints listed in (2), which could be inefficient in practice.
+We can now convert (PSDP-$\mathbb{C}$}) to a real SDP:
+(PSDP-$\mathbb{R}$}) $\begin{cases}\sup\limits_{Y\in{\mathbf{S}}^{2n}}&\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(H_{R})+{\mathscr{A}}_{I}(H_{I})=b_{R},\\ &{\mathscr{A}}_{R}(H_{I})-{\mathscr{A}}_{I}(H_{R})=b_{I},\\ &Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0.\end{cases}$
+As mentioned in the introduction, to feed the PSD constraint in (PSDP-$\mathbb{R}$}) to an SDP solver, we need to include also the extra $n(n+1)$ affine constraints listed in (2), which could be inefficient in practice.
 Below we show that by taking a dual point of view, we can actually get rid of this issue.
-Before formulating the dual problem of ($-{\mathbb{C}}$), we explicitly give the adjoint operator of ${\mathscr{A}}$.
+Before formulating the dual problem of (PSDP-$\mathbb{C}$}), we explicitly give the adjoint operator of $\mathscr{A}$}.
 Lemma 2.1.
@@
 which yields ${\mathscr{A}}^{*}(y)={\mathcal{H}}\left(\sum_{i=1}^{m}y_{i}A_{i}\right)$.
-Then by the convex duality theory, the dual problem of ($-{\mathbb{C}}$) reads as
-($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y\in{\mathbb{C}}^{m}}&\langle b,y\rangle_{{\mathbb{R}}}\\ \,\,\,\rm{s.t.}&{\mathscr{A}}^{*}(y)\succeq C.\end{cases}$
+Then by the convex duality theory, the dual problem of (PSDP-$\mathbb{C}$}) reads as
+(DSDP-$\mathbb{C}$}) $\begin{cases}\inf\limits_{y\in{\mathbb{C}}^{m}}&\langle b,y\rangle_{{\mathbb{R}}}\\ \,\,\,\rm{s.t.}&{\mathscr{A}}^{*}(y)\succeq C.\end{cases}$
 Assume $y=y_{R}+y_{I}\hbox{\bf{i}}$ with $y_{R},y_{I}\in{\mathbb{R}}^{m}$. Using Lemma 2.1, we deduce that ${\mathscr{A}}^{*}(y)=U+V\hbox{\bf{i}}$ with
@@
 $V\coloneqq\frac{1}{2}\sum_{i=1}^{m}{\mathcal{I}}(y_{i}){\mathcal{R}}(A_{i}-A_{i}^{\intercal})+{\mathcal{R}}(y_{i}){\mathcal{I}}(A_{i}-A_{i}^{\intercal}).$
-Thus, we can convert ($-{\mathbb{C}}$) to a real SDP by using the equivalent condition (1):
-($-{\mathbb{R}}$) $\begin{cases}\inf\limits_{y_{R},y_{I}\in{\mathbb{R}}^{m}}&b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}\\ \,\,\,\,\,\,\,\,\rm{s.t.}&\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\succeq 0.\end{cases}$
-Let $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]\in{\mathbf{S}}^{2n}$ be the dual PSD variable of ($-{\mathbb{R}}$) with $X_{1},X_{2},X_{3}\in{\mathbb{R}}^{n\times n}$.
-Then the Lagrangian associated with ($-{\mathbb{R}}$) given by
+Thus, we can convert (DSDP-$\mathbb{C}$}) to a real SDP by using the equivalent condition (1):
+(DSDP-$\mathbb{R}$}) $\begin{cases}\inf\limits_{y_{R},y_{I}\in{\mathbb{R}}^{m}}&b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}\\ \,\,\,\,\,\,\,\,\rm{s.t.}&\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\succeq 0.\end{cases}$
+Let $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]\in{\mathbf{S}}^{2n}$ be the PSD variable of (DSDP-$\mathbb{R}$}) with $X_{1},X_{2},X_{3}\in{\mathbb{R}}^{n\times n}$.
+Then the Lagrangian associated with (DSDP-$\mathbb{R}$}) is given by
 $\,L(X,y_{R},y_{I})$
@@
 $\,+\langle b_{I}-{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})+{\mathscr{A}}_{I}(X_{1}+X_{2}),y_{I}\rangle.$
-Therefore, the dual problem of ($-{\mathbb{R}}$) can be written down as
-($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{X\in{\mathbf{S}}^{2n}}&\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(X_{1}+X_{2})+{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal})=b_{R},\\ &{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})-{\mathscr{A}}_{I}(X_{1}+X_{2})=b_{I},\\ &X=\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix}\succeq 0.\end{cases}$
+Therefore, the dual problem of (DSDP-$\mathbb{R}$}) can be written down as
+(PSDP-$\mathbb{R}$}’) $\begin{cases}\sup\limits_{X\in{\mathbf{S}}^{2n}}&\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(X_{1}+X_{2})+{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal})=b_{R},\\ &{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})-{\mathscr{A}}_{I}(X_{1}+X_{2})=b_{I},\\ &X=\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix}\succeq 0.\end{cases}$
 Theorem 2.3.
-($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{R}}$) (in the sense that they share the same optimum).
-As a result, ($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{C}}$).
-In addition, if $X^{\star}=\left[\begin{smallmatrix}X_{1}^{\star}&(X_{3}^{\star})^{\intercal}\\ X_{3}^{\star}&X_{2}^{\star}\end{smallmatrix}\right]$ is an optimal solution to ($-{\mathbb{R}}$), then $H^{\star}=(X^{\star}_{1}+X_{2}^{\star})+(X^{\star}_{3}-(X_{3}^{\star})^{\intercal})\hbox{\bf{i}}$ is an optimal solution to ($-{\mathbb{C}}$).
+(PSDP-$\mathbb{R}$}) is equivalent to (PSDP-$\mathbb{R}$}) (in the sense that they share the same optimum).
+As a result, (PSDP-$\mathbb{R}$}) is equivalent to (PSDP-$\mathbb{C}$}).
+In addition, if $X^{\star}=\left[\begin{smallmatrix}X_{1}^{\star}&(X_{3}^{\star})^{\intercal}\\ X_{3}^{\star}&X_{2}^{\star}\end{smallmatrix}\right]$ is an optimal solution to (PSDP-$\mathbb{R}$}), then $H^{\star}=(X^{\star}_{1}+X_{2}^{\star})+(X^{\star}_{3}-(X_{3}^{\star})^{\intercal})\hbox{\bf{i}}$ is an optimal solution to (PSDP-$\mathbb{C}$}).
 Proof 2.4.
-Let us denote the optima of ($-{\mathbb{R}}$) and ($-{\mathbb{R}}$) by $v$ and $v^{\prime}$, respectively.
-Suppose that $Y=\left[\begin{smallmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$).
-Then one can easily check that $X\coloneqq\frac{1}{2}Y$ is a feasible solution to ($-{\mathbb{R}}$).
+Let us denote the optima of (PSDP-$\mathbb{R}$}) and (PSDP-$\mathbb{R}$}) by $v$ and $v^{\prime}$, respectively.
+Suppose that $Y=\left[\begin{smallmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{smallmatrix}\right]$ is a feasible solution to (PSDP-$\mathbb{R}$}).
+Then one can easily check that $X\coloneqq\frac{1}{2}Y$ is a feasible solution to (PSDP-$\mathbb{R}$}).
 Moreover, we have $\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle=\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle$ and it follows $v\leq v^{\prime}$.
-On the other hand, suppose $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$).
-We then have
+On the other hand, suppose $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]$ is a feasible solution to (PSDP-$\mathbb{R}$}).
+Since PSDness is preserved under basis transformation, we have
 $\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}^{-1}X\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}=\begin{bmatrix}X_{2}&-X_{3}\\ -X_{3}^{\intercal}&X_{1}\end{bmatrix}\succeq 0,$
-and thus
+and since PSDness is preserved under addition, we further have
 $Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\coloneqq\begin{bmatrix}X_{1}+X_{2}&X_{3}^{\intercal}-X_{3}\\ X_{3}-X_{3}^{\intercal}&X_{1}+X_{2}\end{bmatrix}\succeq 0.$
-One can easily see that $Y$ is a feasible solution to ($-{\mathbb{R}}$) and in addition, it holds $\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle=\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$.
+One can easily see that $Y$ is a feasible solution to (PSDP-$\mathbb{R}$}) and in addition, it holds $\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle=\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$.
 Thus $v\geq v^{\prime}$, which proves the equivalence.
 The latter statement of the theorem is clear from the above arguments.
-In contrast to ($-{\mathbb{R}}$), the PSD constraint in ($-{\mathbb{R}}$) is straightforward, and thus no extra affine constraint is required.
-This is why the conversion ($-{\mathbb{R}}$) is more appealing than ($-{\mathbb{R}}$) from the computational perspective.
+In contrast to (PSDP-$\mathbb{R}$}), the PSD constraint in (PSDP-$\mathbb{R}$}) does not entails any special structure, and thus no extra affine constraint is required to describe it.
+This is why the conversion (PSDP-$\mathbb{R}$}) is more appealing than (PSDP-$\mathbb{R}$}) from the computational perspective.
 Remark 2.5.
-A similar reformulation to ($-{\mathbb{R}}$) but for a restricted class of complex SDP relaxations of multiple-input multiple-output detection has appeared in [11].
+A similar reformulation to (PSDP-$\mathbb{R}$}) but for a restricted class of complex SDP relaxations of multiple-input multiple-output detection has appeared in [13].
 3 Application to complex SDP relaxations for CPOPs
-In this section, we apply the reformulation ($-{\mathbb{R}}$) to complex SDP relaxations arising from the complex moment-HSOS hierarchy for CPOPs. A CPOP is given by
+In this section, we apply the reformulation (PSDP-$\mathbb{R}$}) to complex SDP relaxations arising from the complex moment-HSOS hierarchy for CPOPs. A CPOP is given by
 (CPOP) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&f({\mathbf{z}},\overline{{\mathbf{z}}})=\sum_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}\\ \,\,\rm{s.t.}&g_{i}({\mathbf{z}},\overline{{\mathbf{z}}})=\sum_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}\geq 0,\quad i\in[t],\end{cases}$
 where $\overline{{\mathbf{z}}}\coloneqq(\overline{z}_{1},\ldots,\overline{z}_{s})$ stands for the conjugate of complex variables ${\mathbf{z}}\coloneqq(z_{1},\ldots,z_{s})$.
-The functions $f,g_{1},\ldots,g_{t}$ are real-valued polynomials and their coefficients satisfy $b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$, $g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{g^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$.
+The functions $f,g_{1},\ldots,g_{t}$ are real-valued polynomials in $\mathbf{z}$} and $\overline{{\mathbf{z}}}$ (their coefficients satisfy $b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$, $g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{g^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$).
 The support of $f$ is defined by $\hbox{\rm{supp}}(f)\coloneqq\{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\mid b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$.
@@
 Fix a $d\in{\mathbb{N}}$.
-Let $y=(y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}})_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}}\subseteq{\mathbb{C}}$ be a sequence indexed by $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}$ and satisfying $y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{y_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$.
+Let $y=(y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}})_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}}$ be a complex sequence indexed by $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}$ and satisfying $y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{y_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$.
 Let $L_{y}$ be the linear functional defined by
@@
 The complex moment matrix ${\mathbf{M}}_{d}(y)$ associated with $y$ is the Hermitian matrix indexed by ${\mathbb{N}}^{s}_{d}$ such that
-$[{\mathbf{M}}_{d}(y)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}({\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
+${\mathbf{M}}_{d}(y)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}({\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
 Suppose that $g=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}{\mathbf{z}}^{{\boldsymbol{\beta}}^{\prime}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}^{\prime}}$ is a complex polynomial.
 The complex localizing matrix ${\mathbf{M}}_{d}(gy)$ associated with $g$ and $y$ is the Hermitian matrix indexed by ${\mathbb{N}}^{s}_{d}$ such that
-$[{\mathbf{M}}_{d}(gy)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}(g{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}y_{{\boldsymbol{\beta}}+{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}+{\boldsymbol{\gamma}}^{\prime}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
-Let $d_{0}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ and $d_{i}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ for $i\in[t]$, where ||$\cdot$|| ⋅ | denotes the sum of entries.
+${\mathbf{M}}_{d}(gy)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}(g{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}y_{{\boldsymbol{\beta}}+{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}+{\boldsymbol{\gamma}}^{\prime}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
+Let $d_{0}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ and $d_{i}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ for $i\in[t]$, where $|\cdot|$ denotes the sum of entries.
 Let $d_{\min}\coloneqq\max\,\{d_{0},d_{1},\ldots,d_{m}\}$.
 For any $d\geq d_{\min}$, the $d$-th ($d$ is called the relaxation order) complex moment relaxation for (CPOP) is given by
-($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y}&L_{y}(f)=\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&{\mathbf{M}}_{d}(y)\succeq 0,\\ &{\mathbf{M}}_{d-d_{i}}(g_{i}y)\succeq 0,\quad i\in[t],\\ &y_{\mathbf{0},\mathbf{0}}=1.\end{cases}$
-($-{\mathbb{C}}$) and its dual problem form the complex moment-HSOS hierarchy of (CPOP). For more details on this hierarchy, we refer the reader to [10, 15].
+(Mom-$\mathbb{C}$}) $\begin{cases}\inf\limits_{y}&L_{y}(f)=\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&{\mathbf{M}}_{d}(y)\succeq 0,\\ &{\mathbf{M}}_{d-d_{i}}(g_{i}y)\succeq 0,\quad i\in[t],\\ &y_{\mathbf{0},\mathbf{0}}=1.\end{cases}$
+(Mom-$\mathbb{C}$}) and its dual problem form the complex moment-HSOS hierarchy of (CPOP). For more details on this hierarchy, we refer the reader to [12, 17].
 For any $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d}$, we associate it with a matrix $A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{\omega_{s,d}\times\omega_{s,d}}$ defined by
-(6) $[A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}1,&\text{if }({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
+(6) $A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}1,&\text{if }({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
 Moreover, for each $i\in[t]$, we associate any $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}$ with a matrix $A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{C}}^{\omega_{s,d-d_{i}}\times\omega_{s,d-d_{i}}}$ defined by
-(7) $[A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}},&\text{if }({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
+(7) $A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}},&\text{if }({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
 Now for each $i=0,1,\ldots,t$, we define the $\mathbb{R}$}-linear operator ${\mathscr{A}}^{i}$ by
@@
 ${\mathbf{M}}_{d-d_{i}}(g_{i}y)=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}}A_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}^{i}y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=({\mathscr{A}}^{i})^{*}(y),\quad i=0,1,\ldots,t.$
-Therefore, one can rewrite ($-{\mathbb{C}}$) as follows:
-($-{\mathbb{C}}$’) $\begin{cases}\inf\limits_{y}&\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&({\mathscr{A}}^{i})^{*}(y)\succeq 0,\quad i=0,1,\ldots,t,\\ &y_{\mathbf{0},\mathbf{0}}=1,\end{cases}$
+Therefore, one can rewrite (Mom-$\mathbb{C}$}) as follows:
+(Mom-$\mathbb{C}$}’) $\begin{cases}\inf\limits_{y}&\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&({\mathscr{A}}^{i})^{*}(y)\succeq 0,\quad i=0,1,\ldots,t,\\ &y_{\mathbf{0},\mathbf{0}}=1,\end{cases}$
 whose dual reads as
-($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
-Note that we have used the Kronecker delta $\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}$ in ($-{\mathbb{C}}$).
-Let us from now on fix any order “$<$” on ${\mathbb{N}}^{s}$.
+(HSOS-$\mathbb{C}$}) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
+Note that we have used the Kronecker delta $\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}$ in (HSOS-$\mathbb{C}$}).
+Let us from now on fix any order “<<” on ${\mathbb{N}}^{s}$.
 Proposition 3.1.
-($-{\mathbb{C}}$) is equivalent to the following complex SDP:
-($-{\mathbb{C}}$’) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
+(HSOS-$\mathbb{C}$}) is equivalent to the following complex SDP:
+(HSOS-$\mathbb{C}$}’) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
 Proof 3.2.
@@
 $=\sum_{i=0}^{t}\langle\overline{A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}},H^{i}\rangle=\sum_{i=0}^{t}\overline{[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}}.$
-Let ${\mathscr{A}}^{i}_{R},{\mathscr{A}}^{i}_{I}$ be the real linear operator associated with ${\mathscr{A}}^{i}$ which are defined in a similar way as (4) and (5).
-With $H^{i}=H^{i}_{R}+H^{i}_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$, ($-{\mathbb{C}}$) is equivalent to the following real SDP by using the equivalent condition (1):
-($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{\lambda,Y^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &Y^{i}=\begin{bmatrix}H^{i}_{R}&-H^{i}_{I}\\ H^{i}_{I}&H^{i}_{R}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
-On the other hand, by invoking Theorem 2.3, we obtain another equivalent real SDP conversion of ($-{\mathbb{C}}$):
+Let ${\mathscr{A}}^{i}_{R}$ and ${\mathscr{A}}^{i}_{I}$ be the real linear operators associated with ${\mathscr{A}}^{i}$ which are defined in a similar way as (4) and (5).
+With $H^{i}=H^{i}_{R}+H^{i}_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$, (HSOS-$\mathbb{C}$}) is equivalent to the following real SDP by using the equivalent condition (1):
+(HSOS-$\mathbb{R}$}) $\begin{cases}\sup\limits_{\lambda,Y^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &Y^{i}=\begin{bmatrix}H^{i}_{R}&-H^{i}_{I}\\ H^{i}_{I}&H^{i}_{R}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
+On the other hand, by invoking Theorem 2.3, we obtain another equivalent real SDP conversion of (HSOS-$\mathbb{C}$}):
 (8) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
@@
 (8) is equivalent to the following real SDP:
-($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad{\boldsymbol{\beta}}\neq{\boldsymbol{\gamma}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
+(HSOS-$\mathbb{R}$}’) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad{\boldsymbol{\beta}}\neq{\boldsymbol{\gamma}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
 Proof 3.4.
@@
 $=$ $\,\frac{1}{2}\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{R}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})\left([X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}+[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}-[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}-[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}\right)$
-[…]
-where we have used fact that $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$ and $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}$.
-It follows that $[{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\rangle=0$.
+$=$  $\,0,$
+where we have used fact that $(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$ and $(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}$.
+It follows that ${\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\rangle=0$.
 In addition, for each $i=0,1,\ldots,t$ and ${\boldsymbol{\beta}}\in{\mathbb{N}}^{s}_{d}$, we have
@@
 $=$ $\,\frac{1}{2}\sum_{\begin{subarray}{c}({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})\in{\mathbb{N}}^{s}_{d-d_{i}}\times{\mathbb{N}}^{s}_{d-d_{i}}\\ ({\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime})\in\hbox{\rm{supp}}(g)\\ ({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\beta}})\end{subarray}}{\mathcal{I}}\left(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}}\right)\left([X^{i}_{1}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}+[X^{i}_{2}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}\right)$
-[…]
-where we have used fact that ${\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}})={\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+\overline{g}^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})=0$ and $X^{i}_{1},X^{i}_{2}$ are symmetric.
-It follows that $[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\rangle=0$.
+$=$  $\,0,$
+where we have used fact that ${\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}})={\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+\overline{g}^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})=0$ and that $X^{i}_{1},X^{i}_{2}$ are symmetric.
+It follows that ${\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\rangle=0$.
 Putting all above together yields (9).
 We have proved the following theorem.
+Theorem 3.5.
 Before closing the section, we compare complexity of different real SDP reformulations for complex SDP relaxations of (CPOP) in Table 1.
@@
 $n_{\rm{sdp}}$: the maximal size of SDP matrix, $m_{\rm{sdp}}$: the number of affine constraints.
-($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
+(HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$})
 $n_{\rm{sdp}}$ $2\omega_{s,d}$ $2\omega_{s,d}$
@@
 For comparison, we also include the results of directly solving complex SDPs obtained with Hypatia 0.8.1 [8] in Sections 4.14.4.
-All numerical experiments were performed on a desktop computer with Intel(R) Core(TM) i9-10900 CPU@2.80GHz and 64G RAM.
-When presenting the results, the column labelled by ‘opt’ records the optimum and the column labelled by ‘time’ records running time in seconds.
-Moreover, the symbol ‘-’ means the SDP solver runs out of memory, and the symbol ‘*’ means running time exceeds 10000 seconds.
+We emphasis that the purpose of numerical experiments is to compare different formulations of complex SDPs and a comparison between solvers is not our focus in this paper.
+All numerical experiments were performed on a desktop computer with an Intel(R) Core(TM) i9-10900 CPU@2.80GHz and 64G RAM.
+When presenting the results, the column labelled by ‘opt’ records the optimum and the column labelled by ‘time’ records computational time in seconds.
+Moreover, the symbol ‘-’ means the SDP solver runs out of memory, and the symbol ‘*’ means computational time exceeds 10000 seconds.
+4.1 The complex matrix completion problem
+The complex matrix completion problem seeks to recover a low-rank matrix $M\in{\mathbb{C}}^{s\times t}$ from a subset of entries $\{M_{ij}\}_{(i,j)\in\Omega}$ [6, 7], which could be tackled via the convex optimization problem:
+(10) $\begin{cases}\inf\limits_{Z\in{\mathbb{C}}^{s\times t}}&\|Z\|_{*}\\ \,\,\,\,\,\,\rm{s.t.}&Z_{ij}=M_{ij},\quad\forall(i,j)\in\Omega,\end{cases}$
+where $\|Z\|_{*}\coloneqq\hbox{\rm{Tr}}(Z^{{\mathrm{H}}}Z)^{\frac{1}{2}}$ is the nuclear norm of $Z$. One can equivalently cast (10) as a complex SDP:
+(11) $\begin{cases}\inf\limits_{X\in{\mathrm{H}}_{n}}&\hbox{\rm{Tr}}(X)\\ \,\,\,\rm{s.t.}&Z_{ij}=M_{ij},\quad\forall(i,j)\in\Omega,\\ &X=\begin{bmatrix}U&Z^{{\mathrm{H}}}\\ Z&V\end{bmatrix}\succeq 0.\end{cases}$
+In this subsection, we generate random instances of the complex matrix completion problem (10) as follows: 1) select $\Omega\subseteq[s]\times[s]$ uniformly at random from all subsets with cardinality $10s$ for $s\in\{20,30,40,50,60,70\}$; 2) set $M=M_{1}M_{2}^{{\mathrm{H}}}$ with the real and imaginary parts of entries of $M_{1}\in{\mathbb{C}}^{s\times 5}$ and $M_{2}\in{\mathbb{C}}^{s\times 5}$ being selected i.i.d.
+from the standard normal distribution.
+The related results are shown in Table 2.
+From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is $1\sim 3$ magnitudes faster than the reformulation (HSOS-$\mathbb{R}$}).
+Table 2: The complex matrix completion problem.
+$s$ (PSDP-$\mathbb{R}$}) (PSDP-$\mathbb{R}$}) (PSDP-$\mathbb{C}$})
+$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
+20 1948 315.882 0.62 308 315.882 0.06 315.882 3.46
+30 4168 456.591 5.38 508 456.591 0.11 456.591 31.1
+40 7160 670.611 62.7 680 670.611 0.37 670.611 153
+50 11010 861.510 175 910 861.510 0.77 861.510 626
+60 15602 925.590 573 1082 925.590 1.07 925.590 1899
+70 21056 1108.93 1337 1316 1108.93 1.81 1108.93 4870
 4.2 Minimizing a random complex quartic polynomial over the unit sphere
-Our first example is to minimize a complex quartic polynomial over the unit sphere:
+Our second example is to minimize a complex quartic polynomial over the unit sphere:
 (12) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&[{\mathbf{z}}]_{2}^{{\mathrm{H}}}Q[{\mathbf{z}}]_{2}\\ \,\,\,\rm{s.t.}&|z_{1}|^{2}+\cdots+|z_{s}|^{2}=1,\end{cases}$
-where $[{\mathbf{z}}]_{2}$ is the column vector of monomials in ${\mathbf{z}}$ up to degree two and $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the standard normal distribution.
+where ${\mathbf{z}}]_{2}$ is the column vector of monomials in $\mathbf{z}$} up to degree two and $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the standard normal distribution.
 We approach (12) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations.
 The related results are shown in Table 3.
-From the table, we see that the reformulation ($-{\mathbb{R}}$) is several ($2\sim 7$) times as fast as the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows.
-On the other hand, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
+From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is several ($2\sim 7$) times as fast as the reformulation (HSOS-$\mathbb{R}$}), and the speedup becomes more significant as the SDP size grows.
 Table 3: Minimizing a random complex quartic polynomial over the unit sphere.
-$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
+$s$ $d$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})
 $m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
@@
 4.3 Minimizing a random complex quartic polynomial with unit-norm variables
-The second example is to minimize a random complex quartic polynomial with unit-norm variables:
+The next example is to minimize a random complex quartic polynomial with unit-norm variables:
 (13) $\begin{cases}\inf\limits_{{\mathbf{z}}\in{\mathbb{C}}^{s}}&[{\mathbf{z}}]_{2}^{{\mathrm{H}}}Q[{\mathbf{z}}]_{2}\\ \,\,\,\rm{s.t.}&|z_{i}|^{2}=1,\quad i=1,\ldots,s,\end{cases}$
@@
 The related results are shown in Table 4.
-From the table, we see that the reformulation ($-{\mathbb{R}}$) is about one magnitude faster than the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows.
-Again, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
+From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is about one magnitude faster than the reformulation (HSOS-$\mathbb{R}$}), and again the speedup becomes more significant as the SDP size grows.
 Table 4: Minimizing a random complex quartic polynomial with unit-norm variables.
-$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
+$s$ $d$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})
 $m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
@@
 The results are displayed in Table 5.
-From the table we see that the reformulation ($-{\mathbb{R}}$) is $1.5\sim 2$ times as fast as the reformulation ($-{\mathbb{R}}$).
-Moreover, for this problem, solving the original complex SDP with Hypatia is extremely slow probably due to the fact that the SDP contains many PSD blocks.
+From the table we see that the reformulation (HSOS-$\mathbb{R}$}) is $1.5\sim 2$ times as fast as the reformulation (HSOS-$\mathbb{R}$}).
+Moreover, for this problem, solving the original complex SDP with Hypatia is extremely slow probably due to the fact that the SDPs contain many PSD blocks.
 Table 5: Minimizing a randomly generated sparse complex quartic polynomial over multi-spheres.
-$l$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
+$l$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})
 $m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
@@
 Mathematically, it can be formulated as the following CPOP:
-(13) $\begin{cases}\inf\limits_{V_{i},S_{k}^{g}}&\sum_{k\in G}\left(\mathbf{c}_{2k}({\mathcal{R}}(S_{k}^{g}))^{2}+\mathbf{c}_{1k}{\mathcal{R}}(S_{k}^{g})+\mathbf{c}_{0k}\right)\\ \,\,\,\textrm{s.t.}&\angle V_{r}=0,\\ &\mathbf{S}_{k}^{gl}\leq S_{k}^{g}\leq\mathbf{S}_{k}^{gu},\quad\forall k\in G,\\ &\boldsymbol{\upsilon}_{i}^{l}\leq|V_{i}|\leq\boldsymbol{\upsilon}_{i}^{u},\quad\forall i\in N,\\ &\sum_{k\in G_{i}}S_{k}^{g}-\mathbf{S}_{i}^{d}-\mathbf{Y}_{i}^{sh}|V_{i}|^{2}=\sum_{(i,j)\in E_{i}\cup E_{i}^{R}}S_{ij},\quad\forall i\in N,\\ &S_{ij}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})\frac{|V_{i}|^{2}}{|\mathbf{T}_{ij}|^{2}}-\overline{\mathbf{Y}}_{ij}\frac{V_{i}\overline{V}_{j}}{\mathbf{T}_{ij}},\quad\forall(i,j)\in E,\\ &S_{ji}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})|V_{j}|^{2}-\overline{\mathbf{Y}}_{ij}\frac{\overline{V}_{i}V_{j}}{\overline{\mathbf{T}}_{ij}},\quad\forall(i,j)\in E,\\ &|S_{ij}|\leq\mathbf{s}_{ij}^{u},\quad\forall(i,j)\in E\cup E^{R},\\ &\boldsymbol{\theta}_{ij}^{\Delta l}\leq\angle(V_{i}\overline{V}_{j})\leq\boldsymbol{\theta}_{ij}^{\Delta u},\quad\forall(i,j)\in E,\\ \end{cases}$
 […]
+where $V_{i}$ is the voltage, $S_{k}^{g}$ is the power generation, $S_{ij}$ is the power flow (all are complex variables; $\angle\cdot$ stands for the angle of a complex number) and all symbols in boldface are constants.
+Notice that $G$ is the collection of generators and $N$ is the collection of buses.
+For a full description on the AC-OPF problem, we refer the reader to [3] as well as [5].
 We select test cases from the AC-OPF library PGLiB-OPF [3].
@@
 The results are displayed in Table 6.
-From the table, we again see that the reformulation ($-{\mathbb{R}}$) is several ($1.4\sim 5$) times as fast as the reformulation ($-{\mathbb{R}}$).
+From the table, we again see that the reformulation (HSOS-$\mathbb{R}$}) is several ($1.4\sim 5$) times as fast as the reformulation (HSOS-$\mathbb{R}$}).
 Table 6: The results for the AC-OPF problem. $s$: the number of CPOP variables; $t$: the number of CPOP constraints.
-Case $s$ $t$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
+Case $s$ $t$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$})
 $m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+14_ieee 19 147 2346 $1.9940\text{e}3$ 0.19 422 $1.9940\text{e}3$ 0.10
+30_ieee 36 297 4828 $8.1959\text{e}3$ 0.73 836 $8.1960\text{e}3$ 0.37
+30_as 36 297 4828 $5.0371\text{e}2$ 0.55 836 $5.0371\text{e}2$ 0.24
+39_epri 49 361 5270 $1.3568\text{e}5$ 0.74 966 $1.3579\text{e}5$ 0.54
+89_pegase 101 1221 57888 $9.4098\text{e}4$ 63.6 10262 $9.4101\text{e}4$ 15.1
+57_ieee 64 563 11102 $3.6644\text{e}4$ 2.36 2008 $3.6644\text{e}4$ 1.06
+118_ieee 172 1325 25374 $9.3216\text{e}4$ 8.27 4471 $9.3216\text{e}4$ 2.68
+162_ieee_dtc 174 1809 64874 $1.0492\text{e}5$ 43.4 11327 $1.0495\text{e}5$ 13.8
+179_goc 208 1827 25712 $6.0859\text{e}5$ 10.3 4368 $6.0860\text{e}5$ 3.57
+240_pserc 383 3039 52172 $2.8153\text{e}6$ 31.9 9243 $2.8170\text{e}6$ 10.7
+300_ieee 369 2983 53946 $5.3037\text{e}5$ 40.6 9647 $5.3037\text{e}5$ 10.6
+500_goc 671 5255 90502 $3.9697\text{e}5$ 89.8 15918 $3.9697\text{e}5$ 25.4
+588_sdet 683 5287 79362 $1.9799\text{e}5$ 91.7 13933 $1.9749\text{e}5$ 21.3
+793_goc 890 7019 104978 $1.1194\text{e}5$ 105 18536 $1.1222\text{e}5$ 31.5
+1888_rte 2178 18257 280580 $1.2537\text{e}6$ 939 47205 $1.2545\text{e}6$ 180
+2000_goc 2238 23009 455530 $9.1876\text{e}5$ 2087 77974 $9.1881\text{e}5$ 439
+Summary of the benchmark.
+The benchmark confirms our theoretical findings and demonstrates that our new real reformulation of complex SDPs is indeed more efficient than the classical real reformulation.
+Especially, the speedup factor gets even bigger as the SDP size grows and may reach several magnitudes on certain instances.
+It should be also noted that solving the original complex SDP with Hypatia is even slower than solving the classical real reformulation with MOSEK, probably because MOSEK is commercial and the implementations of SDP algorithms based on real numbers are more mature and robust.
 Acknowledgments
 The authors would like to thank Jurij Volčič for helpful comments on an earlier preprint of this note.
+Funding
+This work was jointly funded by National Key R&D Program of China under grant No.
+2023YFA1009401 & 2022YFA1005102, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No.
+12201618 & 12171324.
+Conflict of interest
+The authors declare that they have no conflict of interest.
+Data availability
+The author confirms that all data generated or analysed during this study are included in this article.